OpenStudy (anonymous):

A square pyramid has a slant height of 17 inches and a height of 15 inches. What is its surface area if the square base has sides of length 16 inches?

OpenStudy (mertsj):

Base = 16^2-256 4 triangular faces = 4 (1/2)(16)(17)=544 Total of all 5 faces = 256 + 544=800

OpenStudy (mertsj):